Simplify the following fractions
7 over 10 plus 3 over 8

Answer by Alan3354(69443)

7/10 + 3/8
--------------
Find a common denominator.
The product always works: 8*10 = 80
--> 56/80 + 30/80 = 86/80
= 43/40